OpenStudy (anonymous):

I NEED HELP I WILL GIVE ANYTHING MEDALS WHATEVERR YOU WANT I NEED HELP NOW!!!

OpenStudy (anonymous):

Preston is making cookies. He has 4 over 5 cup of flour in a bag. He then used two 1 over 10 cups of flour for his recipe. Which of the following best explains the amount of flour left in the bag?

OpenStudy (anonymous):

6 over 10, because 4 over 5 minus 2 times 1 over 10 equals 4 over 5 minus 3 over 10 5 over 10, because 4 over 5 minus 2 times 1 over 10 equals 4 over 5 minus 3 over 10 6 over 10, because 4 over 5 minus 2 times 1 over 10 equals 4 over 5 minus 2 over 10 5 over 10, because 4 over 5 minus 2 times 1 over 10 equals 4 over 5 minus 2 over 10

OpenStudy (anonymous):

@iGreen PLEASE ANYONE SAVE ME!

OpenStudy (ahvei):

i think its B

OpenStudy (igreen):

\(\sf\dfrac{1}{10} + \dfrac{1}{10} =~ \! ?\)

OpenStudy (anonymous):

2/10?

OpenStudy (rockinhood):

1/5 I green

OpenStudy (rockinhood):

xD

OpenStudy (igreen):

Yep! Now divide the numerator and denominator by 2.

OpenStudy (igreen):

2 / 2 = ? 10 / 2 = ?

OpenStudy (anonymous):

1 and 5

OpenStudy (igreen):

Oh, wait..nevermind..convert 4/5 into a denominator of 10. 4 * 2 = ? 5 * 2 = ?

OpenStudy (anonymous):

(-_-) i dont know how to do that cas im dumb

OpenStudy (igreen):

Multiply those two.. * is multiplication

OpenStudy (anonymous):

4*2=8 5*2=10

OpenStudy (anonymous):

please hurry

OpenStudy (igreen):

Correct, so we have 8/10 cups of flour in the bag. Now what's \(\sf\dfrac{8}{10}-\dfrac{2}{10}\)?

OpenStudy (anonymous):

-0.12

OpenStudy (anonymous):

6/10

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

or am i seriously retarded

OpenStudy (igreen):

Yep, 6/10 is correct.

OpenStudy (anonymous):

ok

OpenStudy (igreen):

So which one should be our answer?

OpenStudy (anonymous):

a or c

OpenStudy (igreen):

Yes, and which of those is correct?

OpenStudy (anonymous):

i think its c

OpenStudy (anonymous):

please ?

OpenStudy (igreen):

You are correct, because we have 2/10

OpenStudy (anonymous):

YUSSSS

OpenStudy (anonymous):

one more Q and A?

OpenStudy (igreen):

had*

OpenStudy (anonymous):

Simplify (4.5)(5)(−2). 45 4.5 −4.5 −45

OpenStudy (igreen):

Sure, just close this and open a new one.

OpenStudy (anonymous):

ok one second